# Digital-physical testbed for ship autonomy studies in the Marine Cybernetics Laboratory basin

### v2

@@ 1 Introduction
Autonomous vessels have the potential to significantly influence the maritime sector by enhancing safety, operational efficiency, and economic viability (gu2021autonomous; munim2019autonomous). Recent advances in sensor technology, artificial intelligence, and marine cybernetics have spurred the development of prototype autonomous vessels of various sizes (schiaretti2017survey). Additionally, large-scale initiatives have emerged to demonstrate the real-world feasibility of uncrewed ships under diverse conditions (reddy2019zero). These efforts are largely driven by benefits such as increased safety (through reduced human error), lower operating costs, and reduced emissions, all of which hold significant promise for reshaping shipping operations and logistics (gu2021autonomous; psaraftis2023logistics).
To bridge the gap between concept and full-scale deployment, researchers and industry leaders have developed or repurposed vessels of varying sizes as experimental platforms for autonomous navigation (schiaretti2017survey). The milliAmpere1 (hinostroza2024_OMAE2025) and milliAmpere2 (eide2025autonomous) ferries—both built at half-scale for urban waterways—have facilitated the development and validation of key autonomy functions such as sensor fusion, collision avoidance, and automated docking under real harbor conditions (martinsen2019autonomous). Yara Birkeland (Skredderberget2025) is a fully electric container ship providing a full-scale test platform for uncrewed cargo operations. Furthermore, the crewed research vessel R/V Gunnerus (marley2023four) is equipped with conventional ship systems and an extensive suite of sensors and data acquisition tools, and it provides a platform for advanced scientific studies. In the commercial scale, companies like Maritime Robotics and SeaRobotics each offer a fleet of autonomous surface vessels (ASVs) of multiple sizes for robotics applications and hydrographic surveys, while Reach Subsea, Ocean Infinity, and Fugro each develop their fleets of autonomous vessels specifically designed for offshore operations.
@@
In this work, we introduce a digital-physical testbed for Maritime Autonomous Surface Ships (MASS) that combines high-fidelity simulations, reduced-order models, and real-time sensor feedback within a unified framework. Our testbed facilitates a smooth transition from simulation to physical experimentation, supported by a digital twin environment in Unity for visualization and remote control. This setup aims to enhance the development and validation of autonomy algorithms by providing a cost-effective and flexible platform. The primary contribution is an open, modular software-hardware stack—featuring curated datasets and a validated hydrodynamic model library—that transforms the Marine Cybernetics Laboratory basin into a benchmark platform for ship autonomy research.
@@ 2.1 Overview
Cybership I pettersen1998underactuated End of life
Cybership II skjetne2005maneuvering End of life
Cybership III smogeli2006control End of life
C/S Enterprise I skaatun2011development Operational
C/S Saucer idland2015marine Operational
C/S Inocean Cat I Drillship bjorno2016thruster Operational
C/S Voyager elvenes2023development [^1] C/S Jonny has been renamed to C/S Voyager. Operational
R/V milliAmpere1 brekke2022milliampere Operational
R/V milliAmpere2 eide2025autonomous Operational
R/V Gunnerus marley2023four Operational
@@ 3.1.1 High-fidelity simulation model
The high-fidelity simulation framework mcsimpy [^2] The mcsimpy open-source can be accessed at https://github.com/NTNU-MCS/mcsimpy, written in Python, replicates the dynamics of the C/S fleet by incorporating wave-excitation models and realistic hydrodynamic coefficients. Specifically, it uses JONSWAP (hasselmann1973measurements) and Pierson–Moskowitz (pierson1964proposed) wave spectra to generate diverse sea states, while hydrodynamic parameters are derived from WAMIT (wamit7v5) and VERES (shipx). mcsimpy supports both 3 and 6 degrees of freedom (DOF) models, making it suitable for various vessel configurations, including the R/V Gunnerus. Built on the studies by mo2023real,hygen2023deterministic as well as the Marine Systems Simulator (perez2006overview), this framework provides a versatile platform for simulating control and navigation algorithms under realistic environmental conditions. The intended use of the simulation environment is for control algorithm development, controller tuning, and serving as a fast, flexible sandbox for machine learning (ML) and reinforcement learning (RL) applications.
@@ 3.1.2 Reduced-order model simulation
We employ a lightweight simulation environment [^3] The simulator for the simplified hydrodynamic model is open-source and available at https://github.com/NTNU-MCS/shoeboxpy that approximates vessel dynamics using reduced-order models. This environment is primarily used for rapid testing of software interfaces and high-level control logic, as it shares the same programming interface as the physical platform.
@@
where the Coriolis matrices $\mathbf{C}_{\mathrm{RB}}$ and $\mathbf{C}_{\mathrm{A}}$ incorporate rigid-body and added-mass terms, respectively, $\tau$ represents loads acting on the body, and $\mathbf{J}$ is the transformation matrix between the body and inertial frames (fossen2021handbook). For small roll and pitch angles, a diagonal restoring term $\mathbf{g}(\eta)$ provides simple restoring forces, heave, roll, and pitch. Numerically, we integrate $\eta$ and $\nu$ using a fourth-order Runge–Kutta step at each timestep $\Delta t$. Although this reduced-order approach omits detailed thruster and environmental modeling, it offers sufficient fidelity for early-stage verification and greatly reduces computational overhead.
@@ 3.1.3 Simplified dynamics with sensor models
Stonefish (cieslak2019stonefish) is a high performance and real-time capable robotics simulator written in C++. Here, it is used for providing a sandbox for image processing, supports a wide range of sensor types (such as cameras, inertial measurement units, and sonars), and has simplified collision and wave dynamics. Thanks to its flexibility, Stonefish is ideal for evaluating algorithms like localization, object detection, and platform interaction that requires environmental variability and realistic sensor feedback. Stonefish bridges the gap between reduced-order simulation models and large-scale real-world tests by simulating important elements of the environment, and allows validation of control and perception algorithms.
@@ 3.2 Digital Twin
A digital twin environment has been developed to support testing and visualization of vessel behavior in a virtual representation of the MC-Lab  (premraj2023development). It consists of Unity-implemented 3D models of the MC-Lab and selected vessels from the C/S fleet, including the C/S Enterprise I, C/S Saucer, and C/S Inocean Cat I Drillship. This setup enables the monitoring and demonstration of control algorithms in a more intuitive format and can be used to support testing in early development phases.
The digital twin system is paired with a remote control center, which allows operators to control the physical vessels in the MC-Lab from a separate location (halvorsen2024combining). Command can be performed manually using joysticks or Metaquest 2 headset and hand controls [^4] Link to the video demonstration: https://www.youtube.com/watch?v=fqEdp8VXAqE, as well as using these interfaces to provide guidance commands like waypoints and speed references for automatic control. Vessel state information, such as position and orientation, is transmitted over the site-to-site network to provide remote monitoring capabilities. While some limitations exist due to hardware constraints—particularly in the responsiveness of the monitoring—the system has been tested in both simulation and physical experiments and has been shown to function reliably under the available conditions.
@@ 3.3.1 Core modules
All the vessels in the C/S fleet operate on a shared software stack [^5] C/S software suite can be accessed at https://github.com/NTNU-MCS/cybership_software_suite. The software is developed in a monolithic architecture, with all components maintained in a single repository. This structure simplifies version control, integration, and testing, and is supported by a continuous integration and deployment (CI/CD) pipeline that automates the build and deployment process for each vessel in the testbed.
The core software is written in Python and C++, combining the flexibility and rapid development capabilities of Python with the performance and low-level control afforded by C++. Robot Operating System 2  (macenski2022robot) (ROS 2) serves as the middleware layer, enabling efficient inter-process communication and seamless integration with additional components and external programs. This middleware facilitates modularity, allowing various control, sensing, and estimation modules to operate independently while maintaining synchronized communication across the system.
@@ 3.3.2 Control and estimation modules
Each vessel comes with a set of default software modules, allowing for immediate testing and validation of standard behaviors. These include a thrust allocation module for distribution of forces to the thrusters, a single-input-single-output (SISO) proportional-integrated-derivative (PID) based feedback velocity controller, third-order reference filter integrated with a proportional-derivative (PD) pose controller, and finally an extended Kalman filter (EKF) from the robot_localization package (MooreStouchKeneralizedEkf2014).
@@
We employ the maneuvering-based dynamic thrust allocation method (gezer2024maneuvering), which formulates the thrust allocation task as a dynamic control problem. This approach uses a control Lyapunov function (CLF) and a control barrier function (CBF) to handle rate limitations and force saturations, ensuring smooth and feasible thrust commands.
@@
with $\mathbf{\Omega}=\mathrm{diag}(\omega_{1},\omega_{2},\omega_{3})$ controlling the filter bandwidth, and $\mathbf{\Delta}=\mathrm{diag}(\delta_{1},\delta_{2},\delta_{3})$ controlling the filter damping (fossen2021handbook). Other types of reference models may also be implemented; see (fernandes2012reference). A forward-Euler discretization step is sufficient to update the filter state, providing smooth desired pose, velocity, and acceleration outputs for the DP controller. The default included controller is simply a PD control law to generate the desired thrust loads for the vessel to closely follow the reference trajectory using
@@
The 4-corner test (skjetne2017amos) was conducted using the pose controller, and the results in Fig. 4 show the performance differences between simulation and physical implementation. The simulation achieved a position RMSE of 0.062m, yaw RMSE of $0.439^{\circ}$, and velocity RMSE of 0.031m/s, whereas the physical vessel showed 0.180m, $4.295^{\circ}$, and 0.036m/s, respectively.
@@ 3.4 Available data
To build realistic simulators of the C/S fleet, we have generated and publicly released comprehensive hydrodynamic datasets [^6] Dataset is accessible through https://github.com/NTNU-MCS/mcmodels. We performed stationkeeping hydrodynamic analyses using WAMIT (wamit7v5); a boundary-element software grounded in potential flow theory. The submerged portions of each hull were meshed from approximate 3D scans with some geometric simplifications. To optimize computation, we simulated only half of each hull by exploiting ship symmetry. From these runs, we obtained hydrostatic coefficients, linear hydrodynamic loads, second-order mean drift forces and moments, and response amplitude operators (RAOs) for the six rigid-body motions. Linear hydrodynamic loads (added mass and damping) were derived from the radiation problem, while wave excitations were computed from the diffraction problem over headings ranging from $0^{\circ}$ to $180^{\circ}$. The RAOs were based on vessel mass alone, excluding mooring or viscous effects.
@@ 4 Physical infrastructure
The MC-Lab provides a versatile and well-instrumented environment for testing model-scale marine vessels in realistic wave conditions. Its primary facility is a wave basin measuring $32m$ from the flap-type wave maker to the beach and 6.4m wide, with a maximum depth of $1.5m$. The wave maker flap spans the full basin width, and is capable of generating waves up to $0.25m$ in amplitude at controlled frequencies, and a wave-damping beach is located at the far end that minimizes reflections for cleaner experiments. This allows users to produce both regular and irregular wave spectra, such as JONSWAP, in either manual or automated modes. Conductive and ultrasonic probes can be placed throughout the basin to monitor wave characteristics. A Motion Capture system offers real-time object tracking (to within $1cm$ in position and $0.5^{\circ}$ in orientation). Additional technical specifications and usage details can be found on the web pages of MC-lab. Within 2030, the MC-Lab will be succeeded by a new model basin for academic research, called the Hydro-Cybernetics Laboratory in the House of Archimedes as part of the new Norwegian Ocean Technology Centre (NTNU_Norwegian_Ocean_Technology_Centre). This more modern facility will be a larger test basin, approximately $65m$ in length, $8m$ in width, and $3m$ in depth, further expanding research capabilities in hydrodynamics, marine cybernetics, and ocean technology.
@@ 5 Conclusions
We plan to expand the simulation capabilities by incorporating forward-speed hydrodynamic data from ShipX (shipx), extending high-fidelity simulators to cover relevant speed ranges for the C/S fleet, and publishing experimental datasets for benchmarking and reproducibility.

### v4

@@ 1 Introduction
Autonomous vessels have the potential to significantly influence the maritime sector by enhancing safety, operational efficiency, and economic viability [10, 21]. Recent advances in sensor technology, artificial intelligence, and marine cybernetics have spurred the development of prototype autonomous vessels of various sizes [29]. Additionally, large-scale initiatives have emerged to demonstrate the real-world feasibility of uncrewed ships under diverse conditions [28]. These efforts are largely driven by benefits such as increased safety (through reduced human error), lower operating costs, and reduced emissions, all of which hold significant promise for reshaping shipping operations and logistics [10, 27].
To bridge the gap between concept and full-scale deployment, researchers and industry leaders have developed or repurposed vessels of varying sizes as experimental platforms for autonomous navigation [29]. The milliAmpere1 [13] and milliAmpere2 [5] ferries—both built at half-scale for urban waterways—have facilitated the development and validation of key autonomy functions such as sensor fusion, collision avoidance, and automated docking under real harbor conditions [18]. Yara Birkeland [34] is a fully electric container ship providing a full-scale test platform for uncrewed cargo operations. Furthermore, the crewed research vessel R/V Gunnerus [17] is equipped with conventional ship systems and an extensive suite of sensors and data acquisition tools, and it provides a platform for advanced scientific studies. In the commercial scale, companies like Maritime Robotics and SeaRobotics each offer a fleet of autonomous surface vessels (ASVs) of multiple sizes for robotics applications and hydrographic surveys, while Reach Subsea, Ocean Infinity, and Fugro each develop their fleets of autonomous vessels specifically designed for offshore operations.
@@
In this work, we introduce a digital-physical testbed for Maritime Autonomous Surface Ships (MASS) that combines high-fidelity simulations, reduced-order models, and real-time sensor feedback within a unified framework. Our testbed facilitates a smooth transition from simulation to physical experimentation, supported by a digital twin environment in Unity for visualization and remote control. This setup aims to enhance the development and validation of autonomy algorithms by providing a cost-effective and flexible platform. The primary contribution is an open, modular software-hardware stack—featuring curated datasets and a hydrodynamic model library—that transforms the Marine Cybernetics Laboratory basin into a benchmark platform for ship autonomy research.
@@ 2.1 Overview
Cybership I [24] End of life
Cybership II [32] End of life
Cybership III [35] End of life
C/S Enterprise I [31] Operational
C/S Saucer [15] Operational
C/S Inocean Cat I Drillship [1] Operational
C/S Voyager [6] [^1] C/S Jonny has been renamed to C/S Voyager. Operational
R/V milliAmpere1 [2] Operational
R/V milliAmpere2 [5] Operational
R/V Gunnerus [17] Operational
@@ 3.1.1 High-fidelity simulation model
The high-fidelity simulation framework mcsimpy [^2] The mcsimpy open-source can be accessed at https://doi.org/10.5281/zenodo.17274093, written in Python, replicates the dynamics of the C/S fleet by incorporating wave-excitation models and realistic hydrodynamic coefficients. Specifically, it uses JONSWAP [12] and Pierson–Moskowitz [25] wave spectra to generate diverse sea states, while hydrodynamic parameters are derived from WAMIT [36] and VERES [30]. mcsimpy supports both 3 and 6 degrees of freedom (DOF) models, making it suitable for various vessel configurations, including the R/V Gunnerus. Built on the studies by [19, 14] as well as the Marine Systems Simulator [23], this framework provides a versatile platform for simulating control and navigation algorithms under realistic environmental conditions. The intended use of the simulation environment is for control algorithm development, controller tuning, and serving as a fast, flexible sandbox for machine learning (ML) and reinforcement learning (RL) applications.
@@ 3.1.2 Reduced-order model simulation
We employ a lightweight simulation environment [^3] The simulator for the simplified hydrodynamic model is open-source and available at https://doi.org/10.5281/zenodo.17277252 that approximates vessel dynamics using reduced-order models. This environment is primarily used for rapid testing of software interfaces and high-level control logic, as it shares the same programming interface as the physical platform.
@@
where the Coriolis matrices $\mathbf{C}_{\mathrm{RB}}$ and $\mathbf{C}_{\mathrm{A}}$ incorporate rigid-body and added-mass terms, respectively, $\tau$ represents loads acting on the body, and $\mathbf{J}$ is the transformation matrix between the body and inertial frames [8]. For small roll and pitch angles, a diagonal restoring term $\mathbf{g}(\eta)$ provides simple restoring forces in heave, roll, and pitch. Numerically, we integrate $\eta$ and $\nu$ using a fourth-order Runge–Kutta step at each timestep $\Delta t$. Although this reduced-order approach omits detailed thruster and environmental modeling, it offers sufficient fidelity for early-stage verification and greatly reduces computational overhead.
@@ 3.1.3 Simplified dynamics with sensor models
Stonefish [3] is a high performance and real-time capable robotics simulator written in C++. Here, it is used for providing a sandbox for image processing, supports a wide range of sensor types (such as cameras, inertial measurement units, and sonars), and has simplified collision and wave dynamics. Thanks to its flexibility, Stonefish is ideal for evaluating algorithms like localization, object detection, and platform interaction that requires environmental variability and realistic sensor feedback. Stonefish bridges the gap between reduced-order simulation models and large-scale real-world tests by simulating important elements of the environment, and allows validation of control and perception algorithms.
@@ 3.2 Digital Twin
A digital twin environment has been developed to support testing and visualization of vessel behavior in a virtual representation of the MC-Lab  [26]. It consists of Unity-implemented 3D models of the MC-Lab and selected vessels from the C/S fleet, including the C/S Enterprise I, C/S Saucer, and C/S Inocean Cat I Drillship. This setup enables the monitoring and demonstration of control algorithms in a more intuitive format and can be used to support testing in early development phases.
The digital twin system is paired with a remote control center, which allows operators to control the physical vessels in the MC-Lab from a separate location [11]. Command can be performed manually using joysticks or Metaquest 2 headset and hand controls [^4] Link to the video demonstration: https://www.youtube.com/watch?v=fqEdp8VXAqE, as well as using these interfaces to provide guidance commands like waypoints and speed references for automatic control. Vessel state information, such as position and orientation, is transmitted over the site-to-site network to provide remote monitoring capabilities. While some limitations exist due to hardware constraints—particularly in the responsiveness of the monitoring—the system has been tested in both simulation and physical experiments and has been shown to function reliably under the available conditions.
@@ 3.3.1 Core modules
All the vessels in the C/S fleet operate on a shared software stack [^5] C/S software suite can be accessed at https://doi.org/10.5281/zenodo.17233653. The software is developed in a monolithic architecture, with all components maintained in a single repository. This structure simplifies version control, integration, and testing, and is supported by a continuous integration and deployment (CI/CD) pipeline that automates the build and deployment process for each vessel in the testbed.
The core software is written in Python and C++, combining the flexibility and rapid development capabilities of Python with the performance and low-level control afforded by C++. Robot Operating System 2  [16] (ROS 2) serves as the middleware layer, enabling efficient inter-process communication and seamless integration with additional components and external programs. This middleware facilitates modularity, allowing various control, sensing, and estimation modules to operate independently while maintaining synchronized communication across the system.
@@ 3.3.2 Control and estimation modules
Each vessel comes with a set of default software modules, allowing for immediate testing and validation of standard behaviors. These include a thrust allocation module for distribution of forces to the thrusters, a single-input-single-output (SISO) proportional-integrated-derivative (PID) based feedback velocity controller, third-order reference filter integrated with a proportional-derivative (PD) pose controller, and finally an extended Kalman filter (EKF) from the robot_localization package [20].
@@
We employ the maneuvering-based dynamic thrust allocation method [9], which formulates the thrust allocation task as a dynamic control problem. This approach uses a control Lyapunov function (CLF) and a control barrier function (CBF) to handle rate limitations and force saturations, ensuring smooth and feasible thrust commands.
@@
with $\mathbf{\Omega}=\mathrm{diag}(\omega_{1},\omega_{2},\omega_{3})$ controlling the filter bandwidth, and $\mathbf{\Delta}=\mathrm{diag}(\delta_{1},\delta_{2},\delta_{3})$ controlling the filter damping [8]. Other types of reference models may also be implemented; see [7]. A forward-Euler discretization step is sufficient to update the filter state, providing smooth desired pose, velocity, and acceleration outputs for the DP controller. The default included controller is simply a PD control law to generate the desired thrust loads for the vessel to closely follow the reference trajectory using
@@
The 4-corner test [33] was conducted using the pose controller, and the results in Fig. 4 show the performance differences between simulation and physical implementation. The simulation achieved a position RMSE of 0.062m, yaw RMSE of $0.439^{\circ}$, and velocity RMSE of 0.031m/s, whereas the physical vessel showed 0.180m, $4.295^{\circ}$, and 0.036m/s, respectively.
@@ 3.4 Available data
To build realistic simulators of the C/S fleet, we have generated and publicly released comprehensive hydrodynamic datasets [^6] Dataset is accessible through https://doi.org/10.5281/zenodo.17274087. We performed stationkeeping hydrodynamic analyses using WAMIT [36]; a boundary-element software grounded in potential flow theory. The submerged portions of each hull were meshed from approximate 3D scans with some geometric simplifications. To optimize computation, we simulated only half of each hull by exploiting ship symmetry. From these runs, we obtained hydrostatic coefficients, linear hydrodynamic loads, second-order mean drift forces and moments, and response amplitude operators (RAOs) for the six rigid-body motions. Linear hydrodynamic loads (added mass and damping) were derived from the radiation problem, while wave excitations were computed from the diffraction problem over headings ranging from $0^{\circ}$ to $180^{\circ}$. The RAOs were based on vessel mass alone, excluding mooring or viscous effects.
@@ 4 Physical infrastructure
The MC-Lab provides a versatile and well-instrumented environment for testing model-scale marine vessels in realistic wave conditions. Its primary facility is a wave basin measuring $32m$ from the flap-type wave maker to the beach and 6.4m wide, with a maximum depth of $1.5m$. The wave maker flap spans the full basin width, and is capable of generating waves up to $0.25m$ in amplitude at controlled frequencies, and a wave-damping beach is located at the far end that minimizes reflections for cleaner experiments. This allows users to produce both regular and irregular wave spectra, such as JONSWAP, in either manual or automated modes. Conductive and ultrasonic probes can be placed throughout the basin to monitor wave characteristics. A Motion Capture system offers real-time object tracking (to within $1cm$ in position and $0.5^{\circ}$ in orientation). Additional technical specifications and usage details can be found on the web pages of [4]. Within 2030, the MC-Lab will be succeeded by a new model basin for academic research, called the Hydro-Cybernetics Laboratory in the House of Archimedes as part of the new Norwegian Ocean Technology Centre [22]. This more modern facility will be a larger test basin, approximately $65m$ in length, $8m$ in width, and $3m$ in depth, further expanding research capabilities in hydrodynamics, marine cybernetics, and ocean technology.
@@ 5 Conclusions
We plan to expand the simulation capabilities by incorporating forward-speed hydrodynamic data from ShipX [30], extending high-fidelity simulators to cover relevant speed ranges for the C/S fleet, and publishing experimental datasets for benchmarking and reproducibility.
6 Data Availability Statement
The datasets and software that support the findings of this study are openly available. The hydrodynamic datasets used for model development can be accessed at https://doi.org/10.5281/zenodo.17274087. The high-fidelity simulation framework mcsimpy is available at https://doi.org/10.5281/zenodo.17274093, the reduced-order simulator shoeboxpy at https://doi.org/10.5281/zenodo.17277252, and the C/S software suite at https://doi.org/10.5281/zenodo.17233653. All repositories are publicly accessible under open-source licenses.